# A Review of Graph Neural Networks in Epidemic Modeling

### v1

@@ 4.1.3. Intervention Modeling
Intervention modeling offers a detailed perspective on epidemic spread by simulating the behaviors and interactions of individuals within a network based on intervention strategies. This method provides an intricate view of individual actions, mobility patterns, and the likelihood of disease transmission. When combined with GNN, this approach enhances the model’s capability to represent the diversity and complexity inherent in real-world social networks, augmenting the efficacy of intervention strategies. Song et al. (song2020reinforced) introduced a reinforcement learning framework that dynamically optimizes public health interventions to strike a balance between controlling the epidemic and minimizing economic impacts, to reduce infection rates while maintaining economic activities. To delve deeper into the individual underlying dynamics, Meirom et al. (meirom2021controlling) proposed a dual GNN module strategy. One module updates the node representations according to dynamic processes, while the other manages the propagation of long-range information. Subsequently, they employ RL to modulate the dynamics of social interaction graphs and perform intervention actions on them. This approach aims to indirectly curb epidemic spread by strategically altering network structures, thus avoiding direct intervention in the disease process.
IDRLECA (feng2023contact) embodies a novel integration, combining an infection probability model with an innovative GNN design. The infection probability model calculates the current likelihood of each individual’s infection status. This information, along with personal health and movement data, is utilized to forecast virus transmission through human contacts using the GNN:
@@ 4.2.1. Parameter Estimation for Mechanistic Models
This line of research highlights that hybrid models, which integrate neural networks, dynamically adjust the parameters of mechanistic models. This combination enables the analysis of real-time data, thus informing and refining mechanistic models to ensure their simulations accurately mirror the dynamics of actual epidemics (jhun2021effective; tang2023enhancing; lagatta2021epidemiological). Notably, studies like (lagatta2021epidemiological; xie2022visualization) employ GNNs to estimate the contact (transmission) rate, $\beta$, and to monitor the epidemic evolution through the implementation of the SIR model. Further, research (gao2021stan; tang2023enhancing) estimates both the transmission rates $\beta$ and recovery rates $\gamma$ by leveraging outputs from the GNN. This methodology initiates with the utilization of GRU to derive node embeddings h, which subsequently facilitate the calculation of parameters:
@@
To go beyond single-region recognition, MepoGNN (cao2023metapopulation; cao2023mepognn) extends the SIR model to the metapopulation variant (Data-centric_arxiv22; balcan2009multiscale; mousavi2012enhanced), accommodating heterogeneity within populations and incorporating human mobility to model the spread between sub-populations:
@@
Recent work (liu2023epidemiologyaware) integrates the Cola-GNN (deng2020colagnn) framework with the SIR model through the development of Epi-Cola-GNN, introducing a mobility matrix $\Pi$ to capture the dynamics of infectious disease spread across different locations. Within this matrix, $\pi_{ij}$ quantifies the intensity of human mobility from location $i$ to location $j$, offering a nuanced perspective on the spatial transmission of diseases. This incorporation leads to a modification in the representation of infectious cases within the SIR model framework:
@@
Furthermore, they introduce the concept of the Next-Generation Matrix (NGM) (diekmann2010construction), which provides a clearer epidemiological interpretation and more effectively supports both intra-location spread and inter-location transmission influenced by human mobility.
Instead of simply estimating the rate indicator, EpiGCN (han2023devila) innovatively employs three distinct linear layers to transform the node feature into the SIR state, enhancing the model awareness of the available data:
@@ 4.2.2. Mechanistic-Informed Neural Models
Unlike previous methods wherein neural networks dynamically adjust the parameters of mechanistic models based on data inputs, mechanistic-informed neural models utilize domain knowledge from mechanistic models to inform the architecture and learning processes of GNNs. This strategy flexibility allows for a swift adaptation to changing conditions, markedly improving the accuracy of forecasts and the effectiveness of interventions. Certain studies (meznar2021prediction; wang2023wdcip; tomy2022estimating) utilize the SIR model to generate target data by simulating epidemic spreads from individual nodes, which are then employed to train GNNs for downstream tasks. In the context of source detection tasks, such as those discussed in (ru2023inferring; shah2020finding; sha2021source), one-hot encoded node states $x^{t}_{i}\in\{0,1\}^{M}$, with $M$ representing the number of possible states, are used as inputs for the GNN, where the states are defined as either {S, E, I, R} or {S, I, R}. Song et al. (song2020reinforced) integrated SIHR (a variant of SIR) (kermack1927contribution) simulation environment with the RL framework, providing a dynamic model of epidemic progression for the RL agent. This capability allows the agent to account for individuals who are hospitalized, enabling the dynamic modification of mobility control policies.
To explicitly capture causal dynamics, CausalGNN (wang2022causalgnn) introduces a novel approach to causal modeling by leveraging causal features $\textbf{Q}_{t}=(q_{i,t})\in\mathbb{R}^{N\times 4}$, with $q_{i,t}:S_{i}(t),I_{i}(t),R_{i}(t),D_{i}(t)$ representing the cumulative number of individuals in each state of the SIRD model. A causal encoder is then designed to transform these causal features into node embeddings, operating as follows:
@@ 5.1. Epidemic at Scales
Multi-scale data are crucial in epidemiology because they offer comprehensive insights into both intra-region and inter-region relationships, thus aiding in the more accurate modeling of disease spread. Presently, several studies have acknowledged this importance and initiated the integration of multi-scale data into their frameworks  (tangEnhancingSpatialSpread2023; wangWDCIPSpatiotemporalAIdriven2023; qiuMSGNNMultiscaleSpatiotemporal2023; zheng2021hierst). Although these efforts have yielded models capable of accommodating multi-scale data, existing approaches are limited to processing only two predefined scales, such as county-level and state-level data. Looking ahead, there is growing anticipation for the development of novel models capable of incorporating data across multiple dynamic scales and adaptable to diverse epidemiological tasks.
Meanwhile, scalability must also be considered for numerous reasons: (1) A smaller granularity results in the expansion of graph data. (2) Some tasks require real-time processing  (wangWDCIPSpatiotemporalAIdriven2023). While the number of countries or provinces can be small, the graph for individuals or other necessary parts in epidemic models can be extremely large, e.g., contact information graphs in metropolises, which could make the current methods very time-consuming. Furthermore, the use of multi-scale data and the requirements for real-time processing make the problem even harder.
@@ 5.2. Cross-Modality in Epidemiology
The integration of multi-modal data in epidemiological tasks offers a powerful approach for enhancing our understanding of disease transmission dynamics, improving predictive accuracy, enabling early detection and intervention, conducting comprehensive risk assessments, and fostering interdisciplinary collaboration to address public health challenges more effectively. Data from different modalities can not only serve as augmentations for each other but also compensate for noise from single-modality data. In recent years, some works have successfully incorporated multi-modality in GNNs. Although GNNs are very suitable for information aggregation and handling multi-modality data, there has not been much work exploring the multi-modality of GNNs in an epidemiology setting. Some related works  (linGraphNeuralNetwork2023; hanDevilLandscapesInferring2023) have utilized unstructured data like textual or image data to construct node features. However, there is no cross-modality in their work in terms of node features.
@@ 5.3. Epidemic Diffusion Process
The diffusion process, which is the key component in epidemiological tasks, can be both spatial and temporal. All GNN-based methods discussed above involve information aggregation at one or several time points in a discrete manner. Nevertheless, in the real world, disease spreading is a continuous process, which is incompatible with current methods. To address this problem, Continuous GNNs  (xhonneux2020continuous; chamberlain2021grand; poli2019graph), inspired by Neural ODE  (chen2018neural), can be applied to model the continuous spreading process.
Another problem lies in that both disease spreading and infection take time, and they can happen asynchronously. One related work  (puDynamicAdaptiveSpatio2023) considers different time-space effects and models the effects using the attention mechanism. However, it is still done in a discrete manner, creating gaps in the real-world transmission process.
@@ 5.4. Interventions for Epidemics
In epidemiology, control measures are vital for controlling disease spread and safeguarding public health. They include intervention strategies like vaccination, quarantine, and public health education to limit transmission and minimize outbreaks, ultimately saving lives and reducing the burden on healthcare systems  (bhattacharyya2023public; jamison2007disease; luo2021what). Among the methods mentioned in this paper, most of the research incorporates intervention strategy in agent-based models  (fengContactTracingEpidemic2023; meirom2021controlling; song2020reinforced) or other neural models  (pu2023dynamic). Generally, the interventions in these methods include deleting nodes, altering nodes, and altering edge weights. However, each method only includes one type of intervention, either node-level or edge-level. In the real world, however, interventions can happen at different graph levels and also at different scales. To better model the real situation, multi-level and multi-scale interventions need to be introduced.
@@ 5.5. Generating Explainable Predictions
The study of epidemic modeling not only aims for accurate predictions but also emphasizes interpretability. Ideally, experts will rely on epidemic models’ predictions to make informed decisions. However, relying on a model’s predictions becomes risky if the model cannot provide confidence in its forecasts, given the significant consequences of these decisions. Therefore, interpretability is essential and offers various benefits, including understanding disease dynamics, identifying risk factors, and providing measures of uncertainty. Despite the crucial role of interpretability, neural models investigated thus far have not placed significant emphasis on this aspect, with hybrid models primarily relying on mechanistic models to provide explanations. In recent years, there have been some approaches aimed at providing interpretability for general Graph Neural Networks (dai2021towards; ying2019gnnexplainer; yuan2022explainability; nian2024globally), which also hold potential usefulness in epidemiological settings.
@@ 5.6. Handling Challenges from Epidemic Data
The idea of Data-Centric AI (DCAI) has grown more and more important in recent years  (zha2023data), which inspired people to pay attention not only to modeling but also to processing data itself. In epidemiological tasks, there are also many challenges originating from data, e.g., noise, incomplete data, privacy, etc. Although there have not been many works addressing these challenges in GNN-related epidemiological tasks, we expect future works will try to tackle this problem by proposing model-centric and data-centric methods.
@@
Noisy Data. Rodríguez et al.  (rodriguezDataCentricEpidemicForecasting2022) introduced several sources of data in epidemiology and noise naturally exists in these data. For example, while social media can provide information on epidemic progress, it may also create significant noise by spreading rumors and misinformation. For GNNs in epidemiology, noise can exist both at node-level and edge-level. So far, the denoising mechanisms for GNNs in epidemiology have remained to be studied. Fortunately, there has been a wide range of studies on the robustness of GNNs (jin2020graph; jin2021adversarial; zhu2021deep; liu2021elastic), which may be adapted to epidemic data.
@@
Incomplete Data. In addition to noises, epidemic data can also be incomplete. The data-gathering process is not always perfect and can not guarantee the accurate collection of features for every node. This problem may be mitigated during modeling because the blank features can be replaced by the aggregated neighbor features  (tomy2022estimating) and simulation or interpolation can also be used to infer missing features at some time points. Nevertheless, there is not much work studying the influence of incomplete data in an epidemiology setting while using GNN models.
@@
Privacy Protection. In the real world, epidemic data typically includes sensitive information such as individual health status, location, and potentially identifiable details, which, if mishandled, can lead to severe privacy violations and undermine public trust in health systems  (selgelid2016ethics; caals2017ethics; liuyang2022interpretationa). In contemporary research, methods commonly utilize all available epidemic data that span countries or regions. However, this reliance on large-scale data may heighten awareness among government entities and departments regarding data privacy. Owing to the growing focus on data sensitivity, stringent legislations  (IPR_06; ToFairnessinDataSharing_NEJM16) have been introduced to regulate data collection and utilization. Therefore, future work should take more privacy problems into consideration. Federated Graph Learning (FGL)  (FedGCN_arxiv22; FCCL_CVPR22; FLSurveyandBenchmarkforGenRobFair_arXiv23) emerges as a potent solution to these privacy concerns by leveraging the distributed nature of data without necessitating its central aggregation. This approach aligns with the stringent requirements of data privacy regulations by enabling data to remain at its source, thereby minimizing the risks associated with centralized data storage and processing  (fgssl_IJCAI23; FGGP_aaai2024).
@@ 6. Conclusion
[…]
Complete Taxonomy
Table 1. Summary of epidemiological tasks and Methodology.
 Task Paper Methodology Hybrid Graph Construction
Detection SD-STGCN  (sha2021source) GAT + GRU + SEIR Spatial-Temporal Graph; Static Graph Structure
 (ruInferringPatientZero2023) GCN + SIR Spatial-Temporal Graph; Dynamic Graph Structure
 (shah2020finding) GNN + SIR
Surveillance WDCIP  (wang2023wdcip) GAE Spatial Graph; Static Graph Structure
 (songCOVID19InfectionInference2023) GAT/GCN
 (gouarebDetectionPatientsRisk2023) GCN
 (hanDevilLandscapesInferring2023) GCN + Modified SIR
 (shan2023novel) Graph Fourier Transform
Projection MMCA-GNNA (jhun2021effective) GNN + SIR + RL Spatial-Temporal Graph; Static Graph Structure
DURLECA  (song2020reinforced) GNN + RL
IDRLECA  (fengContactTracingEpidemic2023) GNN + RL Spatial-Temporal Graph; Dynamic Graph Structure
 (meirom2021controlling) GNN + RL
Prediction DGDI  (liu2023human) GCN + Self-Attention Spatial Graph; Static Graph Structure
 (sunDeepDynaForecastPhylogeneticinformedGraph2023) GNN + LSTM
DVGSN  (zhang2023predicting) GNN Temporal-Only Graph; Static Graph Structure
STAN  (gao2021stan) GAT + GRU Spatial-Temporal Graph; Static Graph Structure
MSDNet  (tang2023enhancing) GAT + GRU + SIS
SMPNN  (lin2023grapha) MPNN + Autoregression
ATMGNN  (nguyen2023predicting) MPNN/MGNN + LSTM/Transformer
DASTGN  (puDynamicAdaptiveSpatio2023) GNN + Attention + GRU
MSGNN  (qiu2023msgnna) GCN + N-Beats
STEP  (yuSpatiotemporalGraphLearning2023) GCN + Attention + GRU
 (panagopoulos2021transfer) GNN + LSTM
 (meznar2021prediction) Network Centrality + XGBoost
 (tomy2022estimating) GNN + SEIRD
 (lagatta2021epidemiological) GNN + LSTM + SIRD
GSRNN  (li2019study) GNN + RNN Spatial-Temporal Graph; Static Graph Structure
 (fritzCombiningGraphNeural2022) GNN + SDDR
 (xie2022visualization) GCN + Modified SIR
 (moonGraphBasedDeep2023) K-GNN + GRU
 (sesti2021integrating) GNN + LSTM
 (davahli2021predicting) GNN + LSTM
 (murphy2021deep) GNN
 (kapoor2020examining) GNN + MLP
 (croftForecastingInfectionsSpatiotemporal2023) GAT + GRU
Mepo GNN (cao2023metapopulation; caoMepoGNNMetapopulationEpidemic2022) (TCN + GCN) + Modified SIR Spatial-Temporal Graph; Dynamic Graph Structure
Epi-Cola-GNN  (liu2023epidemiologyaware) Cola-GNN + Modified SIR
CausalGNN  (wang2022causalgnn) GNN + SIRD
ATGCN  (wang2022adaptively) GNN + LSTM
HierST  (zheng2021hierst) GNN + LSTM
RESEAT  (moon2023reseat) GNN + Self-Attention
SAIFlu-Net  (jungSelfAttentionBasedDeepLearning2022) GNN + LSTM
Epi-GNN  (xieEpiGNNExploringSpatial2022) GCN + Attention + RNN
Cola-GNN  (deng2020colagnn) GCN + Attention + RNN
 (cui2021unobservables) GNN + Attention + LSTM
 (mahmud2021human) GNN + RNN
 (guo2021research) STGCN
 (duarte2023time) GCRN/GCLSTM

### v2

@@ 4.1.3. Intervention Modeling
Intervention modeling offers a detailed perspective on epidemic spread by simulating the behaviors and interactions of individuals within a network based on intervention strategies. This method provides an intricate view of individual actions, mobility patterns, and the likelihood of disease transmission. When combined with GNN, this approach enhances the model’s capability to represent the diversity and complexity inherent in real-world social networks, augmenting the efficacy of intervention strategies. Song et al. (song2020reinforced, 30) introduced a reinforcement learning framework that dynamically optimizes public health interventions to strike a balance between controlling the epidemic and minimizing economic impacts, to reduce infection rates while maintaining economic activities. To delve deeper into the individual underlying dynamics, Meirom et al. (meirom2021controlling, 93) proposed a dual GNN module strategy. One module updates the node representations according to dynamic processes, while the other manages the propagation of long-range information. Subsequently, they employ RL to modulate the dynamics of social interaction graphs and perform intervention actions on them. This approach aims to indirectly curb epidemic spread by strategically altering network structures, thus avoiding direct intervention in the disease process.
IDRLECA (feng2023contact, 27) embodies a novel integration, combining an infection probability model with an innovative GNN design. The infection probability model calculates the current likelihood of each individual’s infection status. This information, along with personal health and movement data, is utilized to forecast virus transmission through human contacts using the GNN:
@@ 4.2.1. Parameter Estimation for Mechanistic Models
This line of research highlights that hybrid models, which integrate neural networks, dynamically adjust the parameters of mechanistic models. This combination enables the analysis of real-time data, thus informing and refining mechanistic models to ensure their simulations accurately mirror the dynamics of actual epidemics (jhun2021effective, 59, 63, 104). Notably, studies like (lagatta2021epidemiological, 104, 124) employ GNNs to estimate the contact (transmission) rate, $\beta$, and to monitor the epidemic evolution through the implementation of the SIR model. Further, research (gao2021stan, 25, 63) estimates both the transmission rates $\beta$ and recovery rates $\gamma$ by leveraging outputs from the GNN. This methodology initiates with the utilization of GRU to derive node embeddings h, which subsequently facilitate the calculation of parameters:
@@
To go beyond single-region recognition, MepoGNN (cao2023metapopulation, 70, 125) extends the SIR model to the metapopulation variant (Data-centric_arxiv22, 34, 126, 127), accommodating heterogeneity within populations and incorporating human mobility to model the spread between sub-populations:
@@
Recent work (liu2023epidemiologyaware, 28) integrates the Cola-GNN (deng2020colagnn, 23) framework with the SIR model through the development of Epi-Cola-GNN, introducing a mobility matrix $\Pi$ to capture the dynamics of infectious disease spread across different locations. Within this matrix, $\pi_{ij}$ quantifies the intensity of human mobility from location $i$ to location $j$, offering a nuanced perspective on the spatial transmission of diseases. This incorporation leads to a modification in the representation of infectious cases within the SIR model framework:
@@
Furthermore, they introduce the concept of the Next-Generation Matrix (NGM) (diekmann2010construction, 128), which provides a clearer epidemiological interpretation and more effectively supports both intra-location spread and inter-location transmission influenced by human mobility.
Instead of simply estimating the rate indicator, EpiGCN (han2023devila, 129) innovatively employs three distinct linear layers to transform the node feature into the SIR state, enhancing the model awareness of the available data:
@@ 4.2.2. Mechanistic-Informed Neural Models
Unlike previous methods wherein neural networks dynamically adjust the parameters of mechanistic models based on data inputs, mechanistic-informed neural models utilize domain knowledge from mechanistic models to inform the architecture and learning processes of GNNs. This strategy flexibility allows for a swift adaptation to changing conditions, markedly improving the accuracy of forecasts and the effectiveness of interventions. Certain studies (meznar2021prediction, 96, 58, 101) utilize the SIR model to generate target data by simulating epidemic spreads from individual nodes, which are then employed to train GNNs for downstream tasks. In the context of source detection tasks, such as those discussed in (ru2023inferring, 29, 51, 57), one-hot encoded node states $x^{t}_{i}\in\{0,1\}^{M}$, with $M$ representing the number of possible states, are used as inputs for the GNN, where the states are defined as either {S, E, I, R} or {S, I, R}. Song et al. (song2020reinforced, 30) integrated SIHR (a variant of SIR) (kermack1927contribution, 130) simulation environment with the RL framework, providing a dynamic model of epidemic progression for the RL agent. This capability allows the agent to account for individuals who are hospitalized, enabling the dynamic modification of mobility control policies.
To explicitly capture causal dynamics, CausalGNN (wang2022causalgnn, 24) introduces a novel approach to causal modeling by leveraging causal features $\textbf{Q}_{t}=(q_{i,t})\in\mathbb{R}^{N\times 4}$, with $q_{i,t}:S_{i}(t),I_{i}(t),R_{i}(t),D_{i}(t)$ representing the cumulative number of individuals in each state of the SIRD model. A causal encoder is then designed to transform these causal features into node embeddings, operating as follows:
@@ 5.1. Epidemic at Scales
Multi-scale data are crucial in epidemiology because they offer comprehensive insights into both intra-region and inter-region relationships, thus aiding in the more accurate modeling of disease spread. Presently, several studies have acknowledged this importance and initiated the integration of multi-scale data into their frameworks  (tangEnhancingSpatialSpread2023, 85, 131, 132, 72). Although these efforts have yielded models capable of accommodating multi-scale data, existing approaches are limited to processing only two predefined scales, such as county-level and state-level data. Looking ahead, there is growing anticipation for the development of novel models capable of incorporating data across multiple dynamic scales and adaptable to diverse epidemiological tasks.
Meanwhile, scalability must also be considered for numerous reasons: (1) A smaller granularity results in the expansion of graph data. (2) Some tasks require real-time processing  (wangWDCIPSpatiotemporalAIdriven2023, 131). While the number of countries or provinces can be small, the graph for individuals or other necessary parts in epidemic models can be extremely large, e.g., contact information graphs in metropolises, which could make the current methods very time-consuming. Furthermore, the use of multi-scale data and the requirements for real-time processing make the problem even harder.
@@ 5.2. Cross-Modality in Epidemiology
The integration of multi-modal data in epidemiological tasks offers a powerful approach for enhancing our understanding of disease transmission dynamics, improving predictive accuracy, enabling early detection and intervention, conducting comprehensive risk assessments, and fostering interdisciplinary collaboration to address public health challenges more effectively. Data from different modalities can not only serve as augmentations for each other but also compensate for noise from single-modality data. In recent years, some works have successfully incorporated multi-modality in GNNs. Although GNNs are very suitable for information aggregation and handling multi-modality data, there has not been much work exploring the multi-modality of GNNs in an epidemiology setting. Some related works  (linGraphNeuralNetwork2023, 133, 50) have utilized unstructured data like textual or image data to construct node features. However, there is no cross-modality in their work in terms of node features.
@@ 5.3. Epidemic Diffusion Process
The diffusion process, which is the key component in epidemiological tasks, can be both spatial and temporal. All GNN-based methods discussed above involve information aggregation at one or several time points in a discrete manner. Nevertheless, in the real world, disease spreading is a continuous process, which is incompatible with current methods. To address this problem, Continuous GNNs  (xhonneux2020continuous, 134, 135, 136), inspired by Neural ODE  (chen2018neural, 137), can be applied to model the continuous spreading process.
Another problem lies in that both disease spreading and infection take time, and they can happen asynchronously. One related work  (puDynamicAdaptiveSpatio2023, 66) considers different time-space effects and models the effects using the attention mechanism. However, it is still done in a discrete manner, creating gaps in the real-world transmission process.
@@ 5.4. Interventions for Epidemics
In epidemiology, control measures are vital for controlling disease spread and safeguarding public health. They include intervention strategies like vaccination, quarantine, and public health education to limit transmission and minimize outbreaks, ultimately saving lives and reducing the burden on healthcare systems  (bhattacharyya2023public, 138, 139, 140). Among the methods mentioned in this paper, most of the research incorporates intervention strategy in agent-based models  (fengContactTracingEpidemic2023, 60, 93, 30) or other neural models  (pu2023dynamic, 99). Generally, the interventions in these methods include deleting nodes, altering nodes, and altering edge weights. However, each method only includes one type of intervention, either node-level or edge-level. In the real world, however, interventions can happen at different graph levels and also at different scales. To better model the real situation, multi-level and multi-scale interventions need to be introduced.
@@ 5.5. Generating Explainable Predictions
The study of epidemic modeling not only aims for accurate predictions but also emphasizes interpretability. Ideally, experts will rely on epidemic models’ predictions to make informed decisions. However, relying on a model’s predictions becomes risky if the model cannot provide confidence in its forecasts, given the significant consequences of these decisions. Therefore, interpretability is essential and offers various benefits, including understanding disease dynamics, identifying risk factors, and providing measures of uncertainty. Despite the crucial role of interpretability, neural models investigated thus far have not placed significant emphasis on this aspect, with hybrid models primarily relying on mechanistic models to provide explanations. In recent years, there have been some approaches aimed at providing interpretability for general Graph Neural Networks (dai2021towards, 141, 142, 143, 144), which also hold potential usefulness in epidemiological settings.
@@ 5.6. Handling Challenges from Epidemic Data
The idea of Data-Centric AI (DCAI) has grown more and more important in recent years  (zha2023data, 145), which inspired people to pay attention not only to modeling but also to processing data itself. In epidemiological tasks, there are also many challenges originating from data, e.g., noise, incomplete data, privacy, etc. Although there have not been many works addressing these challenges in GNN-related epidemiological tasks, we expect future works will try to tackle this problem by proposing model-centric and data-centric methods.
@@
Noisy Data. Rodríguez et al.  (rodriguezDataCentricEpidemicForecasting2022, 146) introduced several sources of data in epidemiology and noise naturally exists in these data. For example, while social media can provide information on epidemic progress, it may also create significant noise by spreading rumors and misinformation. For GNNs in epidemiology, noise can exist both at node-level and edge-level. So far, the denoising mechanisms for GNNs in epidemiology have remained to be studied. Fortunately, there has been a wide range of studies on the robustness of GNNs (jin2020graph, 147, 148, 149, 150), which may be adapted to epidemic data.
@@
Incomplete Data. In addition to noises, epidemic data can also be incomplete. The data-gathering process is not always perfect and can not guarantee the accurate collection of features for every node. This problem may be mitigated during modeling because the blank features can be replaced by the aggregated neighbor features  (tomy2022estimating, 101) and simulation or interpolation can also be used to infer missing features at some time points. Nevertheless, there is not much work studying the influence of incomplete data in an epidemiology setting while using GNN models.
@@
Privacy Protection. In the real world, epidemic data typically includes sensitive information such as individual health status, location, and potentially identifiable details, which, if mishandled, can lead to severe privacy violations and undermine public trust in health systems  (selgelid2016ethics, 151, 47, 152). In contemporary research, methods commonly utilize all available epidemic data that span countries or regions. However, this reliance on large-scale data may heighten awareness among government entities and departments regarding data privacy. Owing to the growing focus on data sensitivity, stringent legislations  (IPR_06, 153, 154) have been introduced to regulate data collection and utilization. Therefore, future work should take more privacy problems into consideration. Federated Graph Learning (FGL)  (FedGCN_arxiv22, 155, 156, 157) emerges as a potent solution to these privacy concerns by leveraging the distributed nature of data without necessitating its central aggregation. This approach aligns with the stringent requirements of data privacy regulations by enabling data to remain at its source, thereby minimizing the risks associated with centralized data storage and processing  (fgssl_IJCAI23, 158, 159).